OpenStudy (anonymous):

A cylinder's radius is reduced to 2/5 its original size and the height is quadrupled. How has the volume of the cylinder changed? A. 8/25 B. 16/25 C. 64/25 D. 16/5

OpenStudy (anonymous):

let 5r be the radius and h be the height. \[V1=\pi \left( 5r \right)^2h=25\pi r^2h\] again \[radius=\frac{ 2 }{ 5 }\times 5 r=2r,height=4h\] \[V2=\pi \left( 2r \right)^2\left( 4h \right)=16 \pi r^2 h\] \[find \frac{ V2 }{ V1 }\]

OpenStudy (anonymous):

Thank you so.

OpenStudy (anonymous):

yw